OpenStudy (anonymous):

this may be a brain fart, i thought it was five ??? If g(x) = 3(x - 2), find the value of x if g(x) = 6. 12 2.6 8 4

OpenStudy (anonymous):

3(6-2) 3(4) =12

OpenStudy (anonymous):

but

OpenStudy (anonymous):

g(x) = 6 so i would have to plug in a number that would give me 6

OpenStudy (anonymous):

I though g(x)=6 meant that you plug in 6 for x

OpenStudy (anonymous):

Okay, so your answer is either 12 or 4. 6=3(x-2) Divide by 3 2=x-2 Add 2 4=x

OpenStudy (anonymous):

noo the answer is 4

OpenStudy (anonymous):

Alright, sorry I led you astray

OpenStudy (anonymous):

because (2-4)*3=6

OpenStudy (anonymous):

i got my own answer